The sum of an even integer and TWICE the next greater even integer is 8 more than 4 times the greater integer. Find the smaller integer of the pair.
Answer by RAY100(1637)

Let x= smaller even integer,,,,(x+2) next larger even integer
.
(x) + 2(x+2) = 8+ 4(x+2),,,,given relationship
.
x+2x +4 = 8 + 4x +8
.
-12 = x............(x+2) = -10
.
,
check,,,,(-12) + 2(-10) = 8 + 4(-10)
.
-12 -20 = 8 -40 = -32,,,,,,ok
.
.